OpenStudy (anonymous):

Find coefficient of x^2 and x

OpenStudy (anonymous):

Consider the polynomial P(x)= (x-cos 36)(x-cos 84)(x-cos 156)

OpenStudy (loser66):

it's not hard, just tedious, you can do it by yourself :)

OpenStudy (anonymous):

so there must be an easy way

ganeshie8 (ganeshie8):

use Vieta's formulas

OpenStudy (anonymous):

I don't know them ,

ganeshie8 (ganeshie8):

x^3 + bx^2 + cx + d sum of roots = -b sum of products of two roots = c product of roots = d

OpenStudy (anonymous):

is that called vieta i didn't knew sry

ganeshie8 (ganeshie8):

yes but u don't need to know it, just expand the cubic and see what u get for coefficients

OpenStudy (anonymous):

yeah wait

ganeshie8 (ganeshie8):

\[\large P(x)= (x-\cos 36)(x-\cos 84)(x-\cos 156)\] using vieta's formula, coefficient of x^2 would be \(-(\cos 36 + \cos 84 + \cos 156)\) simplify if possible

OpenStudy (anonymous):

yes getting the same

ganeshie8 (ganeshie8):

and that magically turns out to be 0

OpenStudy (anonymous):

0

ganeshie8 (ganeshie8):

how did u get 0 ?

OpenStudy (anonymous):

I had a sum where i got this as i result and it was 0

OpenStudy (anonymous):

had done

OpenStudy (anonymous):

I can evaluate it though , i am not sure about cos 24 and cos96

ganeshie8 (ganeshie8):

ok the x coefficient would be : \[\cos 36 \cos 84 + \cos 84 \cos 156+ \cos 156\cos 36\]

ganeshie8 (ganeshie8):

i need to think a bit on how to simplify these

ganeshie8 (ganeshie8):

it evaluates to -3/4 somehow

OpenStudy (anonymous):

answer is -3/2